# Exploring Lattice Thermal Conductivity Models via Interpretable Deep Learning to Accelerate Materials Discovery

### v7

@@ IV.1 Kolmogorov-Arnold Networks (KANs)
$B_{i,0}(x)$ =$=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1},\\ 0,&\text{otherwise}.\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x).$ (12)
@@ IV.2 Sure Independence Screening & Sparsifying Operator (SISSO)
SISSO constructs a descriptor vector ${{\mathbf{d}}_{n}}$ using features and operators, and models the target vector $\mathbf{P}$ through a linear combination of $n$-dimensional descriptors. The initial feature ${{\mathbf{\Phi}}_{0}}$ serves as the starting point for constructing descriptors and includes easily obtainable physical parameters. The operator set is defined as:
@@ IV.5 Dataset construction & feature preprocessing
The dataset used in this work is sourced from the aflowlib.org database [85], comprising a total of 5,578 entries. For feature selection, the focus was primarily on characteristics related to vibrational, thermodynamic, and mechanical properties at room temperature (300K), which are theoretically associated with LTC according to first-principles researches [86, 87]. For more details, please refer to Table 2. In the modeling aimed at LTC, the model predicts the logarithm of $\kappa$ (log($\kappa$)) rather than the LTC itself. This is because the logarithmic transformation compresses the target value space, thereby enhancing the performance of the ML models [88]. For deep models such as MLPs and KANs, normalization of features is essential as these models rely on gradient descent for parameter optimization. Normalization enhances numerical stability by ensuring that features with different ranges contribute equally to the gradient computations; otherwise, features with larger values might dominate the optimization process [89]. However, this step is unnecessary for XGBoost [90], as it is based on decision trees. Additionally, for interpretability reasons, feature normalization is often omitted in recent works based on SISSO [26, 28, 91]. In this work, we employed Min-Max Normalization, which is expressed as:
@@
where $x$ is the original feature value, ${{x}_{\min}}$ and ${{x}_{\max}}$ are the minimum and maximum values of the feature, respectively, and ${x}^{\prime}$ is the normalized feature value.
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of LTC, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and LTC are computed using ShengBTE [96]. The force constants are generated with Hiphive [97]. For other structures, specific heat is calculated using Phonopy [98], while LTC is determined through GPUMD [99]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.
@@ Acknowledgements.
Supplementary Information
Supplementary Note 1 KAN’s symbolic model on complete dataset
The simplified symbolic model in the main text Eq.(5), derived from key physical features, is sufficient for an accurate description of LTC. However, for reference, we also provide the symbolic expression fitted by KAN using the complete dataset. It is worth noting that, due to the architecture of KAN, this formula is inevitably more complex. Therefore, we present it solely in its Python representation:
1 log(kappa) =
[…]
Supplementary Note 2 Multi-Layer Perceptrons (MLPs)
The theoretical framework of MLPs is well-established. Therefore, in this work, we emphasize the distinctions between MLPs and KANs. Unlike KANs, MLPs are founded on the universal approximation theorem [100], which can be mathematically expressed as:
$f\left(x\right)\approx\sum\limits_{i=1}^{N}{{{\alpha}_{i}}\sigma\left(\omega_{i}^{\text{T}}x+{{b}_{i}}\right)}$ (Supplementary Eq.1)
where $\alpha$ are the coefficients, $\sigma$ is the activation function (we adopted ReLU [101] in this work), ${{\omega}_{i}}$ and ${{b}_{i}}$ denote the weights and biases, $N$ represents the number of neurons in the hidden layer. The expression of MLPs can be simplified to:
$\text{MLP}\left(\mathbf{x}\right)=\left({{\mathbf{W}}_{L-1}}\circ\sigma\circ{{\mathbf{W}}_{L-1}}\circ\sigma\circ\cdots\circ{{\mathbf{W}}_{1}}\circ\sigma\circ{{\mathbf{W}}_{0}}\right)\mathbf{x}$ (Supplementary Eq.2)
In MLPs, only the weights and biases, which serve as linear parameters, are learnable, while the nonlinear activation functions are fixed as hyperparameters.
Supplementary Note 3 eXtreme Gradient Boost (XGB)
XGB represents a great advancement over the traditional Gradient Boosting Decision Tree (GBDT) algorithm. Its innovations include the incorporation of regularization to enhance generalization, a two-step gradient approximation of the objective function to accelerate computational efficiency, column subsampling to mitigate noise and further bolster generalization, and an advanced method for handling missing values by assigning a default classification direction at tree nodes, which renders it particularly effective for sparse datasets [102]. GBDT algorithms, including XGB, are computationally intensive particularly in tasks with high-dimensional features and large datasets [102, 103].
Supplementary Note 4 Crystal Graph Convolutional Neural Network (CGCNN)
CGCNN [104] directly learns from the connection of atoms in a crystal, representing the crystal as a graph with atoms as nodes and bonds as edges. This graph-based approach enables the model to automatically extract optimal representations for predicting material properties. The key feature of CGCNN is its Convolutional Neural Network (CNN) built on top of these crystal graphs. The convolutional layers update the feature vectors of atoms by incorporating information from neighboring atoms and bonds, using the following graph convolution function:
$v_{i}^{\left(t+1\right)}=\text{Conv}\left(v_{i}^{\left(t\right)},v_{j}^{\left(t\right)},u_{\left(i,j\right)}^{\left(k\right)}\right),{{\left(i,j\right)}_{k}}\in G$ (Supplementary Eq.3)
where $v_{i}^{\left(t\right)}$ is the feature vector of atom i at layer t, and $u_{\left(i,j\right)}^{\left(k\right)}$ is the feature vector of the bond between atoms i and j. After multiple convolutional layers, the pooling layers aggregate the atomic feature vectors into an overall crystal representation, which is then passed to fully connected layers to predict material properties.
Supplementary Note 5 SHAP Indices
Shapley Additive Explanations (SHAP) [105] is derived from the concept of Shapley values [106] in game theory, which were originally employed to assess the contribution of each individual player to the collective gain within a coalition [107]. Shapley value for individual player can be expressed as:
${{\phi}_{i}}=\sum\limits_{S\subseteq N\backslash\left\{i\right\}}{\frac{\left|S\right|!\left(\left|N\right|-\left|S\right|-1\right)!}{\left|N\right|!}}\left(v\left(S\cup\left\{i\right\}\right)-v\left(s\right)\right),i=1,2,\ldots,N$ (Supplementary Eq.4)
where $N$ represents the set of all features, $S$ is any subset of features that does not include feature $i$, $\left|S\right|$ is the number of features in set $S$, is the contribution of feature set $S$ to the model’s prediction output, and $\left(S\cup\left\{i\right\}\right)$ is the contribution of the feature set $S\cup\left\{i\right\}$, which includes feature $i$, to the model’s prediction output. The core concept of SHAP is to generate explanations for individual predictions, which can then be aggregated to provide both local and global insights into the model’s behavior [105], it can be expressed mathematically as:
$\hat{f}\left(\mathbf{x}\right)={{\phi}_{0}}+\sum\limits_{i=1}^{N}{{{\phi}_{i}}\left(\mathbf{x}\right)}$ (Supplementary Eq.5)
where ${{\phi}_{i}}\left(\mathbf{x}\right)$ is the Shapley value of the i-th variable, ${{\phi}_{0}}$ is typically set as the expected value of the training samples, $\hat{f}\left(\mathbf{x}\right)$ represents the original prediction made by the model [108].
Supplementary Note 6 LIME Indices
Local Interpretable Model-agnostic Explanations (LIME) [60] is a perturbation-based method for feature importance analysis that employs a local surrogate model to interpret individual samples. By generating perturbations around the selected sample, a new dataset is created, and an interpretable model (such as linear regression) is trained. The difference between the two models is then measured using an objective function:
$\xi\left(x\right)=\arg\underset{g\in G}{\mathop{\min}}\,L\left(f,g,{{\pi}_{x}}\right)+\Omega\left(g\right)$ (Supplementary Eq.6)
in the expression, $f$ denotes the model to be explained, $g$ represents a simplified model drawn from a set $G$ of potential models (such as linear models), ${{\pi}_{x}}$ quantifies the distance between new samples ${x}^{\prime}$ and the original samples $x$, and $\Omega\left(g\right)$ reflects the complexity of the model $g$.
Supplementary Note 7 Supplementary Tables
Supplementary Table 1: The optimal hyperparameters.
Model Hyperparameter Optimal Value
MLP hidden_sizes [64, 64, 64] (fixed)
learning_rate 0.002587427
batch_size 186
num_epochs 592
XGB n_estimators 489
learning_rate 0.134952344
max_depth 6
subsample 0.824230319
colsample_bytree 0.869698424
reg_alpha 0.9
reg_lambda 37.30538969
KAN width 7
grid 15
k 5
lamb_l1 9.559785284
steps 27
lr 0.47852298
SISSO (10-D) desc_dim 2
nsf 10
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 5
funit (1:10)
nmodels 100
nf_sis 10
SISSO (4-D) desc_dim 2
nsf 4
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 4
funit (1:4)
nmodels 100
nf_sis 10
Supplementary Table 2: Marginal distribution of individual primary features determined by Copula sampling.
Distribution Parameters
$\Theta$ lognorm $\mu=-1.935,\sigma=0.519$
$\gamma$ lognorm $\mu=1.671,\sigma=0.012$
$C_{p}$ lognorm $\mu=-2.611,\sigma=0.778$
$C_{v}$ lognorm $\mu=-2.674,\sigma=0.831$
$\alpha$ lognorm $\mu=-2.370,\sigma=0.507$
$S_{v}$ norm $\mu=0.323,\sigma=0.119$
$F$ lognorm $\mu=-0.027,\sigma=0.102$
$K_{s}$ gamma $k=2.546,\theta=0.102$
$K_{i}$ gamma $k=2.441,\theta=0.104$
Supplementary Note 8 Supplementary Figure
Supplementary Figure 1: Histogram of the marginal distribution of a single key variable.
The variables involved in Table S2 are classified into three types of distributions: Gamma distribution, normal distribution, and log-normal distribution. Their respective probability density functions (PDF) can be expressed as:
$f\left(x\right)$ $=\frac{1}{{\Gamma\left(k\right){\theta^{k}}}}{x^{k-1}}{{\rm{e}}^{-\frac{x}{\theta}}},$ (Supplementary Eq.7)
$f(x)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (Supplementary Eq.8)
$f\left(x\right)$ $=\frac{1}{{x\sigma\sqrt{2\pi}}}{{\rm{e}}^{-\frac{{{{\left[{\ln\left(x\right)-\mu}\right]}^{2}}}}{{2{\sigma^{2}}}}}}.$ (Supplementary Eq.9)

### v8

@@ IV.1 Kolmogorov-Arnold Networks (KANs)
$B_{i,0}(x)$ =$=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1}\\ 0,&\text{otherwise}\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x)$ (11)
@@ IV.2 Sure Independence Screening & Sparsifying Operator (SISSO)
SISSO constructs a descriptor vector ${{\mathbf{d}}_{n}}$ using features and operators, and models the target vector $\mathbf{P}$ through a linear combination of $n$-dimensional descriptors. The vector of primary features, ${{\mathbf{\Phi}}_{0}}$, serves as the starting point for constructing descriptors. The operator set is defined as:
@@ IV.5 Dataset construction & feature preprocessing
The dataset used in this work is sourced from the aflowlib.org database [89], comprising a total of 5,578 entries. For feature selection, the focus was primarily on characteristics related to vibrational, thermodynamic, and mechanical properties at room temperature (300K), which are theoretically associated with $\kappa_{\rm L}$ according to first-principles researches [90, 91]. For more details, please refer to Table 2. In the modeling aimed at $\kappa_{\rm L}$, the model predicts the logarithm of $\kappa$ ($\log(\kappa_{\rm L})$) rather than the $\kappa_{\rm L}$ itself. This is because the logarithmic transformation compresses the target value space, thereby enhancing the performance of the ML models [92]. For deep models such as MLPs and KANs, normalization of features is essential as these models rely on gradient descent for parameter optimization. Normalization enhances numerical stability by ensuring that features with different ranges contribute equally to the gradient computations; otherwise, features with larger values might dominate the optimization process [93]. However, this step is unnecessary for XGBoost [94], as it is based on decision trees. Additionally, for interpretability reasons, feature normalization is often omitted in recent works based on SISSO [26, 28, 95]. In this work, we employed Min-Max Normalization, which is expressed as:
@@
where $x$ is the primary feature value, ${{x}_{\min}}$ and ${{x}_{\max}}$ are the minimum and maximum values of the feature, respectively, and ${x}^{\prime}$ is the normalized feature value.
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of $\kappa_{\rm L}$, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and $\kappa_{\rm L}$ are computed using ShengBTE [100]. The force constants are generated with Hiphive [101]. For other structures, specific heat is calculated using Phonopy [102], while $\kappa_{\rm L}$ is determined through GPUMD [103]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.